# A Backbone for Long-Horizon Robot Task Understanding

### v2

@@ Abstract
End-to-end robot learning, particularly for long-horizon tasks, often results in unpredictable outcomes and poor generalization. To address these challenges, we propose a novel Therblig-based Backbone Framework (TBBF) to enhance robot task understanding and transferability. This framework uses therbligs (basic action elements) as the backbone to decompose high-level robot tasks into elemental robot configurations, which are then integrated with current foundation models to improve task understanding. The approach consists of two stages: offline training and online testing. During the offline training stage, we developed the Meta-RGate SynerFusion (MGSF) network for accurate therblig segmentation across various tasks. In the online testing stage, after a one-shot demonstration of a new task is collected, our MGSF network extracts high-level knowledge, which is then encoded into the image using Action Registration (ActionREG). Additionally, Large Language Model (LLM)-Alignment Policy for Visual Correction (LAP-VC) is employed to ensure precise action execution, facilitating trajectory transfer in novel robot scenarios. Experimental results validate these methods, achieving 94.37% recall in therblig segmentation and success rates of 94.4% and 80% in real-world online robot testing for simple and complex scenarios, respectively. Supplementary material is available at: https://sites.google.com/view/therbligsbasedbackbone/home
@@ I Introduction
A major challenge is data efficiency, as current systems require large, resource-intensive datasets [1, 2, 3]. Models trained for long-horizon tasks excel in simple tasks such as pick&place but struggle with more complex tasks like liquid pouring, which involve multiple steps like grasping, pouring, and releasing. The narrow focus on simple tasks limits these models’ applicability in diverse and complex scenarios. Additionally, end-to-end systems are difficult to explain, analyze, and improve. These systems [4, 5, 6] can learn directly from sensory inputs and discover complex feature representations, eliminating the need for manual feature engineering. However, they are hard to interpret and struggle to transfer skills to new tasks.
To address these challenges, we propose a structured and modular backbone framework. In our context, a backbone refers to a structured framework that decomposes high-level tasks into fundamental units, each representing a specific action or sequence of actions, facilitating better task understanding and transferability [7]. Complex tasks involve rich contact interactions (e.g., multiple points of force application) and operations in dynamic or cluttered environments, requiring the integration of diverse skills and sensor modalities (e.g., vision, force) [8]. Long-horizon tasks require extended sequences of actions, typically involving more than 10 individual steps [9], or over a prolonged period (e.g., several minutes) to achieve a specific goal.
Figure 1: Illustration of the proposed robot task understanding system. By extracting the key backbone of complex tasks and integrating contextual information from one-shot demonstrations, this system efficiently reasons about task-related objects and actions, and generates new trajectories in novel scenarios.
Many models depend heavily on pre-training in specific scenarios [10, 11, 12, 13, 14, 15], limiting their ability to generalize to new environments. Effective scenario generalization is essential for adapting to the complexity and variability of real-world situations [15]. Furthermore, task generalization is crucial, as many systems require extensive examples to learn new tasks, which is impractical in dynamic environments with frequent task variations. Handling a diverse range of tasks is essential for practical robotics. However, existing models often excel only in narrow tasks and struggle with broader activities, reducing their real-world utility. Table I compares various robot learning systems in key capacities, including data efficiency, task complexity, explanation, diversity, generalization, success rate, pre-training scenario requirements, scenario complexity, and multi-modal fusion capability.
@@
Capacity / System Ours Q-attention [12] PerAct [10] Diffuser [13] MimicPlay [14] GLiDE [15] Coarse to Fine [11]
Data Efficiency Highly Efficient Moderately Moderately Efficient Inefficient Highly Efficient Highly Efficient Moderately Efficient
Task Complexity Long-Horizon Short-Horizon Short-Horizon Short-Horizon Long-horizon Short-Horizon Short-Horizon
Task Explanation High Clarity Low Clarity Medium Clarity Low Clarity Medium Clarity High Clarity Medium Clarity
Task Diversity Wide Domain Narrow Domain Narrow Domain Narrow Domain Wide Domain Narrow Domain Wide Domain
Task Generalisation One-shot Few-shot 53 shots 10k shots 20&40 shots Few-shot One-shot
Success Rate 94.4% None *65% *54.4% *55% *99.6% *70%
@@
*The results are from the original presented papers.
To overcome these limitations, we propose a structured and modular backbone framework. This approach enhances the adaptability and efficiency of robot learning systems, improving generalization across various tasks and scenarios while providing a clearer structure for task execution. Our framework (Fig. 1) decomposes robot tasks into two sub-levels: robot trajectory (internal information) and environmental context (external information). The robot trajectory covers the entire motion during the task, while the environmental context includes all objects involved.
By decomposing the robot trajectory, we identify elemental motions, termed ’Therbligs’ (details in Section 2), which serve as backbone elements for understanding the task trajectory and context. Decomposing the environmental context into lower-level elements reveals objects and their attributes, such as mask, position, orientation, and task relevance. As shown in Table I, our approach offers high data efficiency, achieving superior performance with smaller datasets, which is crucial for practical applications. The main contributions of this work are as follows:
  • Therblig-Based Backbone Framework (TBBF): A framework that decomposes complex tasks into fundamental units, enhancing generalization and data efficiency in long-horizon tasks.
  • Meta-RGate SynerFusion Network (MGSF): A network for accurate therblig segmentation during offline training, improving understanding of sequential actions.
  • Action Registration (ActionREG): Integrates therbligs with object configurations for precise action registration and robust task execution.
  • LLM-Alignment Policy for Visual Correction (LAP-VC): Improves tolerance to errors and reduces dependency on highly accurate demonstrations.
We conducted offline and online tests to evaluate the robustness and generalization of our framework. During offline training, the MGSF network segmented therbligs across six tasks: tool-pick&place, crossbeam-cutting, bricks-gluing, tissue-sweeping, surface-wiping, and cup-pouring. In online testing, we collected one-shot demonstrations for five new tasks: board-rolling, foamblock-flipping, plate-scrubbing, spoon-tilting, and paper-stamping.
@@ II Related research
Various intelligent robot systems achieve high accuracy in specific tasks, such as cable routing [16], cloth manipulation [17], and fruit grasping [18]. However, they often struggle to generalize across different tasks. Recently, efforts have been made to develop systems that can handle a variety of tasks [14, 11, 13]. However, these systems typically require large datasets or simple scenarios for generalization and lack a clear backbone for task understanding. To overcome these limitations, we propose using therbligs as the backbone for better task comprehension. Therbligs, which consist of 18 basic motions, provide a systematic way to describe and analyze detailed actions in various tasks [19]. Initially developed for studying human movements, we apply these concepts to visually and conceptually break down the robot’s trajectory during learning and task execution. In robot learning, therbligs offer a clear and structured method to visualize processes, as shown in Fig. 2.
Figure 2: Detailed Decomposition of a Robotic Task into therbligs. The sequence starts with Rest (R), where the robot remains idle. Transport Empty (TE) shows robot moving without any load. Delay (D) represents a pause, allowing for system adjustments or synchronization with other processes. The Grasp (G) phase involves robot’s precise movement to hold the object. In Transport Load (TL), robot carries the object to its intended destination. The Use (U) phase illustrates the robot performing the primary function with the object. Release (R) Load depicts the robot releasing the object.
@@
Figure 3: Overview of the proposed TBBF. This pipeline integrates offline training and online testing stages. During offline training, human experts provide demonstrations and label robot trajectories into therbligs, which are then used to train the MGSF network. In the online testing stage, the trained MGSF network segments new tasks into Therblig-level actions. ActionREG registers these actions into new configurations, and LAP-VC is utilized for error compensation. Finally, YOLOv8 and PCA are used to match new configurations, ensuring robust performance across varied and complex scenarios.
@@
An early contribution in this area is the work by Ahmadzadeh et al. [20], who introduced a novel approach for converting action sequences into symbolic representations. For action modalities, Chen et al. [21] introduced a novel approach that leverages sequential motion primitives through human demonstration. Utilizing a hierarchical BiLSTM classifier, they extracted an intuitive high-level knowledge called therbligs, which reconstructs motion information recorded from human-guided demonstrations, allowing for more general representations for different task decompositions. However, their work remains at a conceptual stage for simple therblig segmentation, and it lacks integration with vision modalities, resulting in limited generalization and scene understanding. For vision modalities, Dessalene et al. [22] introduced a rule-driven, compositional, and hierarchical action modeling method based on therbligs to analyze complex motions. This model features a novel hierarchical architecture comprising a Therblig model and an action model, utilizing vision as a medium for robot action segmentation. However, it lacks integration with action modalities based on human demonstrations, leading to a significant domain gap and inefficient capture of accurate task representations through images alone.
The concurrent study of multimodal approaches (action, vision, language) was introduced by Wang et al. [15]. This research utilized a novel LLM-based system to explore mode families within robot manipulation trajectories. The LLM can generate a list of actions involved in tasks such as scooping marbles from one bowl into another. For instance, the LLM might produce a sequence of verbs such as “reach,” “scoop,” “transport,” and “pour.” This capability allows the system to easily generate goals. These verbs are conceptually close to therbligs, providing a strong basis for task decomposition. However, while this approach closely approximates the therbligs concept, it still does not clearly establish this structure nor utilize it as the backbone for constructing the primary framework. As a result, the system lacks a well-defined structure that could enhance its robustness and reliability in diverse scenarios. Without a clear backbone, the integration of action, vision, and language modalities remains fragmented, leading to potential gaps in task representation and execution efficiency.
@@
Figure 4: Detailed architecture of the MGSF network. The MGSF network integrates BiLSTM and Transformer sub-networks to capture sequential dependencies in robotic tasks. It utilizes a meta-recursive gated fusion mechanism to dynamically combine the outputs of these sub-networks. This fusion process adapts to different tasks, ensuring the effective integration of long-term dependencies.
@@
In our research, we propose to use therbligs as the backbone of a robot intelligent system to enhance task understanding. This therblig-based backbone system represents a significant contribution towards a more structured, interpretable, and adaptable framework for robot task learning and execution. By integrating this approach with the foundation model, we can easily extract the detailed configurations of the objects. Thus, we can create a more robust and flexible model for robotic systems.
@@ III-A TBBF: Explainable robot task understanding framework
In the offline training stage, we utilize the MGSF network to accurately segment tasks into therbligs, providing a detailed breakdown of the task into its constituent motions. During the online testing stage, we collect a one-shot demonstration of a new task, from which the MGSF network extracts high-level knowledge and transforms it into a structured format. This knowledge is encoded into visual data using ActionREG, integrating therbligs with the objects’ configurations in the robot’s visual field to ensure precise action registration. By using therbligs as the backbone, our framework significantly improves data efficiency and task generalization, enabling the robot to handle a wide range of scenarios with robustness and precision. The integration of LAP-VC further ensures that any visual discrepancies are corrected in real time, providing an additional layer of accuracy in task execution. As depicted in Fig. 3, this comprehensive and structured methodology enhances the robot’s ability to adapt to new tasks by leveraging prior knowledge encoded in therbligs, thus improving the interpretability, stability, and transferability of robotic learning systems.
III-B MGSF: Effieicent therbligs segmentation network
@@
  Input: $X\in\mathbb{R}^{n\times d}$, $m\in\mathbb{N}$
  Output: $\hat{Y}\in\mathbb{R}^{n\times O}$
@@
        Update fusion output $F\leftarrow G\odot c_{i}+(1-G)\odot F$where $\odot$ denotes element-wise multiplication}
@@
Our MGSF network, illustrated in Fig. 4 and detailed in Algorithm 1, advances the current state of the art in understanding robot task actions. By combining meta-learning with adaptive gated fusion within a unified framework, this model significantly enhances robots’ ability to comprehend and execute sequential actions across various environments. Inspired by MetaGross [23], our MGSF network incorporates meta-gating and recursive parameterization in a recurrent model. However, MetaGross lacks a dedicated fusion process and struggles to integrate different aspects of the input data effectively, limiting its ability to leverage diverse features and achieve robust performance.
Figure 5: Details of the action registration and context matching process. This figure illustrates the stages of action registration from initial scenarios to new configurations. Key therbligs such as ’Grasp’ and ’Use’ are identified from expert demonstrations using the SAM model and matched to object masks and trajectories. The therblig-based YOLOv8 model is employed for task-related object matching, while inverse kinematics ensures accurate positioning.
To address these limitations, our MGSF model introduces a dynamic hybrid architecture that combines the strengths of both BiLSTM and Transformer sub-networks with a novel adaptive gated fusion mechanism. This architecture features a meta-recursive gated fusion unit that dynamically adapts to integrate model outputs, thereby enhancing performance across diverse tasks. Unlike the static gating in MetaGross, our adaptive gated fusion mechanism allows for more flexible and responsive integration of sequential data, ensuring that long-term dependencies are effectively captured and processed. By leveraging the strengths of both BiLSTM and Transformer sub-networks, the MGSF network excels in handling complex sequences with greater precision. The meta-learning component dynamically adjusts to the changing context of tasks, ensuring that the model remains accurate and applicable across different situations. This fusion process, informed by meta-learning, allows the MGSF model to provide a thorough analysis of each action, thereby improving the robot’s ability to perform a wide range of tasks with enhanced recall and adaptability.

### v3

@@ Abstract
End-to-end robot learning, particularly for long-horizon tasks, often results in unpredictable outcomes and poor generalization. To address these challenges, we propose a novel Therblig-Based Backbone Framework (TBBF) as a fundamental structure to enhance interpretability, data efficiency, and generalization in robotic systems. TBBF utilizes expert demonstrations to enable therblig-level task decomposition, facilitate efficient action-object mapping, and generate adaptive trajectories for new scenarios. The approach consists of two stages: offline training and online testing. During the offline training stage, we developed the Meta-RGate SynerFusion (MGSF) network for accurate therblig segmentation across various tasks. In the online testing stage, after a one-shot demonstration of a new task is collected, our MGSF network extracts high-level knowledge, which is then encoded into the image using Action Registration (ActionREG). Additionally, Large Language Model (LLM)-Alignment Policy for Visual Correction (LAP-VC) is employed to ensure precise action registration, facilitating trajectory transfer in novel robot scenarios. Experimental results validate these methods, achieving 94.37% recall in therblig segmentation and success rates of 94.4% and 80% in real-world online robot testing for simple and complex scenarios, respectively. Supplementary material is available at: https://sites.google.com/view/therbligsbasedbackbone/home
Index Terms:
Deep Learning in Grasping and Manipulation; Manipulation Planning; Learning from Demonstration
@@ I Introduction
Traditional approaches to robot learning typically require large datasets and focus on simpler tasks like pick-and-place operations [1]. These end-to-end methods have shown effectiveness in mapping simple actions to objects within controlled environments [2]. However, when it comes to long-horizon tasks and cluttered scenarios involving multiple steps, multiple unseen objects and intricate object interactions—such as liquid pouring—their effectiveness diminishes. In these situations, actions-objects mapping becomes significantly more challenging. End-to-end systems become inefficient and difficult to train, requiring vast amounts of data and computational resources [3]. Moreover, models trained on specific scenarios often lack the flexibility to adapt to new environments or task variations.
To address these limitations, we propose a novel framework that efficiently reconstructs long-horizon robot tasks by extracting action backbone and registering context information (see Fig. 1). Inspired by the concept of therbligs [4], a set of elemental motions, we introduce the fundamental architecture Therblig-Based Backbone Framework (TBBF). It systematically decomposes complex tasks into fundamental action units—termed ”therbligs.” These therbligs form the foundation for detailed task segmentation, action registration, task-related object reasoning, and trajectory adaptation to new layouts. Thus, our framework offers an interpretable, efficient, and scalable solution for robot task understanding, providing several key benefits:
Figure 1: Concept of the Proposed Robot Task Understanding System: extracts key backbone of complex tasks and uses context from a single demonstration to understand relevant objects and actions.
@@
Capacity / System TBBF(Ours) Q-attention [5] PerAct [6] Diffuser [7] MimicPlay [8] GLiDE [9] Coarse to Fine [10]
Data Efficiency High Moderate Moderate Low High High Moderate
Task Horizon Long-Horizon Short-Horizon Mixed-Horizon Short-Horizon Long-horizon Short-Horizon Short-Horizon
Task Interpretability Well-Structured Unstructured Partially Structured Unstructured Partially Structured Well-Structured Partially Structured
Task Diversity Wide Domain Narrow Domain Wide Domain Narrow Domain Wide Domain Narrow Domain Wide Domain
Task Generalization One-shot Few-shot 53 shots 10k shots 20&40 shots Few-shot One-shot
@@
* Data efficiency evaluates the amount and type of input data required by the system to learn and perform tasks. The Task Horizon indicates whether the system includes long-horizon tasks. Long-horizon tasks require extended sequences of actions, typically involving more than 10 individual steps [11]. Mixed-Horizon refers to tasks that contain both short and long horizons. Task Explanation assesses the clarity and interpretability of the system’s decision-making process, such as action, context information and trajectory transformation. Task diversity assesses type of operation, object diversity, and environmental constraints. Scenario Complexity evaluates the number, diversity, arrangement of objects, and system’s ability to adapt to new layouts (more details in our project website additional materials: https://sites.google.com/view/therbligsbasedbackbone/home).
Interpretability: TBBF enables transparency in the robot learning and execution processes by structuring tasks into distinct therblig sequences. This breakdown supports explainability by allowing each action in a complex task to be represented as interpretable units, making it possible to monitor, analyze, and generalize each action segment clearly and accurately.
Data Efficiency and Robustness: By focusing only on task-relevant actions and objects, TBBF allows the system to disregard irrelevant objects in cluttered scenarios, improving robustness and reducing the need for re-training. This selective attention to critical elements also enables our system to learn from one-shot demonstrations, making it far more data-efficient than models that require extensive datasets for action-object mapping.
Adaptability and Transferability: TBBF excels in generalization, allowing our system to handle new tasks and scenarios with minimal input—only a single image and one trajectory demonstration are required for online testing. By identifying therblig indices associated with specific actions, TBBF can adapt trajectories for new layouts by transferring task-related object changes across different environments. This adaptability is essential for real-world applications where tasks often vary dynamically.
As shown in Table 1, we benchmark TBBF against other architectures, demonstrating its superiority in task diversity, data efficiency, interpretability, and resilience in complex environments. This structured, therblig-based framework fundamentally improves robot learning and automation capabilities, addressing key limitations in existing end-to-end systems. The main contributions of this work include:
  • Therblig-Based Backbone Framework (TBBF): A structured framework that decomposes complex tasks into therbligs, enabling task decomposition, action-object mapping, and trajectory generalization. This framework enhances explainability, generalization, and efficiency in executing complex, long-horizon tasks.
  • Meta-RGate SynerFusion Network (MGSF): A network for precise therblig segmentation, enhancing the understanding of sequential actions in robot tasks.
  • Action Registration (ActionREG): A mechanism that integrates therbligs with object configurations, ensuring accurate action registration and stable task execution.
  • LLM-Alignment Policy for Visual Correction (LAP-VC): A method that leverages large language models for visual error correction, reducing dependency on highly accurate demonstrations and enhancing adaptability.
@@ II Related research
Various intelligent robot systems achieve high accuracy in specific tasks, such as cable routing [12], cloth manipulation [13], and fruit grasping [14]. However, they often struggle to generalize across different tasks. Recently, efforts have been made to develop systems that can handle a variety of tasks [8, 10, 7]. However, these systems typically require large datasets or simple scenarios for generalization and lack a clear backbone for task understanding.
Figure 2: Detailed Decomposition of a Robotic Task into therbligs. The sequence containing: Rest (R), Transport Empty (TE), Delay (D), Grasp (G), Transport Load (TL), Use (U) and Release (R).
@@
Figure 3: Overview of the proposed TBBF. This pipeline integrates offline training and online testing stages. During offline training, human experts provide demonstrations and label robot trajectories into therbligs, which are then used to train the MGSF network. In the online testing stage, the trained MGSF network segments new tasks into Therblig-level actions. ActionREG registers these actions into new configurations, and LAP-VC is utilized for error compensation. Finally, YOLOv8 and PCA are used to match new configurations. Arrows indicate the starting and ending points of the trajectory flow.
@@
Kinematic and dynamic states-based Task Decomposition: Ahmadzadeh et al. [15] introduced a method for converting action sequences into symbolic representations. Building on this, Chen et al. [16] leveraged sequential motion primitives from human demonstrations, using a hierarchical BiLSTM classifier to extract intuitive high-level knowledge called therbligs. This approach enables more general representations for different task decompositions. However, their work remains conceptual, focusing on simple therblig segmentation without integrating vision modalities. This limitation results in reduced generalization and limited scene understanding.
Video-based Task Decomposition: For vision modalities, Dessalene et al. [17] introduced a rule-driven, compositional, and hierarchical action modeling method based on therbligs to analyze complex motions. This model features a novel hierarchical architecture comprising a Therblig model and an action model, utilizing vision as a medium for robot action segmentation. However, it lacks integration with action modalities based on human demonstrations, leading to a significant domain gap and inefficient capture of accurate task representations through images alone.
Language-based Task Decomposition: Large Language Models (LLMs) are utilized in robotics for task decomposition, each with advantages and limitations. Language instructions as input [9][18][19][20][21] enable models like LLaMA and GPT-4 to quickly interpret high-level tasks and generate action sequences without detailed programming, but they lack precise trajectory data for executable paths. Moreover, systems that rely on manually designed action primitives often require extensive human engineering and can struggle to adapt to unseen or highly variable tasks, limiting their scalability. Such primitives may not capture the nuanced spatial or temporal details needed for complex, long-horizon tasks, and transferring them for new layouts.
@@
Figure 4: Detailed architecture of the MGSF network. The MGSF network integrates BiLSTM and Transformer sub-networks to capture sequential dependencies and use a meta-recursive gated fusion mechanism to dynamically combine the outputs of these sub-networks.
@@
In our research, we propose to use therbligs as the backbone of a robot intelligent system to enhance task understanding. This TBBF represents a significant contribution towards a more structured, interpretable, and adaptable framework for robot task learning. By integrating this approach with the foundation model, we can easily extract the detailed configurations of the objects. Thus, we can create a more robust and flexible model for robotic systems.
@@ III-A TBBF: Explainable robot task understanding framework
In the offline training stage, we utilize the MGSF network to accurately segment tasks into therbligs, providing a detailed breakdown of the task into its constituent motions. During the online testing stage, we collect a one-shot demonstration of a new task, from which the MGSF network extracts high-level knowledge and transforms it into a structured format. This knowledge is encoded into visual data using ActionREG, integrating therbligs with the objects’ configurations in the robot’s visual field to ensure precise action registration. By using therbligs as the backbone, our framework significantly improves data efficiency and task generalization, enabling the robot to handle a wide range of scenarios with robustness and precision. The integration of LAP-VC further ensures that any visual discrepancies are corrected in real time, providing an additional layer of accuracy in task execution. As depicted in Fig. 3, this advanced methodology enhances the robot’s ability to adapt to new tasks by leveraging prior knowledge encoded in therbligs, thus improving interpretability, stability, and transferability of robotic learning systems.
Figure 5: Details of the action registration, context matching, and new trajectory generating process. Arrows indicate the direction of trajectory.
III-B MGSF: Efficient therbligs segmentation network
@@
  Input: $X\in\mathbb{R}^{n\times d}$kinematic&dynamic states sequence}
  Output: $\hat{Y}\in\mathbb{R}^{n\times O}$therblig sequence}
@@
        Update fusion output $F\leftarrow G\odot c_{i}+(1-G)\odot F$
@@
Algorithm 1 and Figure 4 Notation: The input $X\in\mathbb{R}^{n\times d}$ represents a sequence of task states (including joint angles, velocity, end-effector position, orientation, force, and torque), where $n$ is the sequence length, and $d$ is the feature dimension. For output, $O\in\mathbb{R}^{n\times k}$ represents the one-hot encoded output, where $n$ is the sequence length and $k$ is the number of classes. Parameters $\theta_{L}$ and $\theta_{T}$ correspond to the BiLSTM and Transformer sub-networks, respectively, each designed to capture different aspects of the task’s sequential dependencies. The meta-parameter vector $M\in\mathbb{R}^{m}$ is initialized to control the gated fusion mechanism dynamically, allowing the model to adapt to task-specific demands. Gate parameters $\theta_{g}$ are generated from $M$ and are used to dynamically control the fusion of BiLSTM and Transformer features, balancing between short-term and long-term dependencies. The recursive fusion step uses $G^{(t)}$ at each time step $t$ to combine features $c_{i}$ and the previous fusion output $F^{(t-1)}$ into the fused feature $F$. The classifier layer, with parameters $\theta_{c}$, uses the final fused feature $F$ to output a predicted sequence of task labels $\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{n}]$, where each $\hat{y}_{i}$ corresponds to a predicted label for each time step in the sequence. The ground truth sequence $Y=[y_{1},y_{2},\dots,y_{n}]$ represents the actual therblig labels (basic action elements) for each time step, with each label encoded in one-hot format. The Binary Cross-Entropy (BCE) loss $\mathcal{L}$ is computed between the predicted sequence $\hat{Y}$ and the ground truth sequence $Y$ to guide the learning process. The Meta Model dynamically updates gate control parameters $W_{g}$ and bias $b_{g}$ for each task via meta parameters $\theta_{m}$ to enhance adaptability across tasks.
Our MGSF network is illustrated in Fig. 4 and detailed in Algorithm 1. By combining meta-learning with adaptive gated fusion within a unified framework, this model significantly enhances robots’ ability to comprehend and execute sequential actions across various environments. Inspired by MetaGross [22], our MGSF network incorporates meta-gating and recursive parameterization in a recurrent model. However, MetaGross lacks a dedicated fusion process and struggles to integrate different aspects of the input data effectively, limiting its ability to leverage diverse features.
To address these limitations, our MGSF model introduces a dynamic hybrid architecture that combines the strengths of both BiLSTM and Transformer sub-networks with a novel adaptive gated fusion mechanism. This architecture features a meta-recursive gated fusion unit that dynamically adapts to integrate model outputs, thereby enhancing performance across diverse tasks. Unlike the static gating in MetaGross, our adaptive gated fusion mechanism allows for more flexible and responsive integration of sequential data, ensuring that long-term dependencies are effectively captured and processed. By leveraging the strengths of both BiLSTM and Transformer sub-networks, the MGSF network excels in handling complex sequences with greater precision. The meta-learning component dynamically adjusts to the changing context of tasks, ensuring that the model remains accurate and applicable across different situations.